# Process-based Inference for Spatial Energetics Using Bayesian Predictive Stacking

### v2

@@ 4. Prediction via stacking
We exploit the analytical closed forms for the posterior distributions and carry out inference using Bayesian stacking (Le and Clarke, 2017; Zhang et al., 2023). In both the discrete-time and continuous-time trajectory models, we are able to obtain closed-form posterior distributions if we fix some hyperparameters in the spatial-temporal covariance structures. We consider a collection of $G$ models, $\{\mathcal{M}_{1},\ldots,\mathcal{M}_{G}\}$, where each $\mathcal{M}_{g}$ is specified by fixing a set of parameters such that the corresponding posterior distribution given dataset $\mathcal{D}$, $p_{g}(\cdot\,|\,\mathcal{D})$, is in closed form.
To be specific, for the discrete time trajectory model in (3), the posterior distribution for $\mathcal{M}_{g}$ is
@@
where $\delta^{2}_{g}=\{\delta^{2}_{\beta,g},\delta^{2}_{z,g}\}$, $\phi_{g}=\{\phi_{1,g},\phi_{2,g}\}$ and $\xi_{g}$ are the fixed values of these parameters for $\mathcal{M}_{g}$. Further, the posterior predictive distributions of $y(\gamma(t_{0}),t_{0})$ and $z(\gamma(t_{0}),t_{0})$ at the new time point $t_{0}$ are calculated from t-distributions, where the details of their arguments are provided in Section 5.3.
@@ 4.2. Predictive stacking of distributions
Spatial energetics seeks full predictive inference on the trajectories entailing interpolation of the latent process at arbitrary points, which subsequently drives predictions for the outcomes. We achieve this by stacking the posterior predictive distributions for each $\mathcal{M}_{g}$ using $\mathcal{D}_{train}$, which is a multivariate t-distribution. Similar to the stacking of means, we consider the weights. Stacking maximizes the score function $S\left(\sum_{g=1}^{G}a_{g}p_{g}(\cdot\mid\mathcal{D}_{train}),q_{t}(\cdot\mid\mathcal{D}_{train})\right)$ to obtain the weights, where $q_{t}$ is a posterior distribution with true underlying parameters. If we employ a logarithmic score, corresponding to the Kullback–Leibler divergence (Yao et al., 2018; Zhang et al., 2023), the weights are obtained as $\operatornamewithlimits{argmax}_{a_{1},\ldots,a_{G}}\sum_{y\in\mathcal{D}_{valid}}\log\left(\sum_{g=1}^{G}a_{g}p_{g}(y\mid\mathcal{D}_{train})\right)$, the logarithm acts on the pseudo-posterior probabilities, given the weighted models. Thus, we define the distributional prediction by maximizing the pseudo-log joint posterior probability. Note that $p_{g}(\cdot\mid\mathcal{D}_{train})$ is a multivariate t-distribution, and hence, evaluating the posterior probability of the validation data is readily available. This optimization problem can be solved as a linearly constrained problem via an adaptive barrier algorithm (Lange, 2010). Algorithm 2 presents the steps involved in stacking of predictive densities.
@@
All the $G$ candidate models in Algorithms 1 and 2 can be computed in parallel and the computation of the weights is negligibly small compared to that of the posterior distribution. Further, the optimization is supported by many packages in various statistical programming languages. In particular, for the subsequent illustrations we employed the (“stats” and “quadprg” packages R Core Team, 2024; Turlach and Weingessel, 2019) in the R statistical computing environment. By contrast, MCMC demands a substantial number of iterations for convergence, and the issue is exacerbated with the larger values of $n$ and $T$.
@@ 5.1. Distribution theory for Bayesian DLMs
Fixing $\delta_{\beta},\delta_{z},\phi$ yields familiar posterior distributions for ${\theta}_{t}$ and $\sigma$, which facilitate stacking. Here, we collect the key recursion equations customarily used in calculating the posterior distribution for (1).
@@
where ${m}_{t,(1:p)},{m}_{t,(p+1:p+n)}$ are the first $p$ elements and the remaining elements of ${m}_{t}$, $W_{t,(1:p,1:p)}$ is the top-left $p\times p$ square of $W_{t}$, $C=(\delta_{z}^{2}K_{\phi}(s,s^{\prime}))_{s,s^{\prime}\in\chi}$, $C_{0}=(\delta_{z}^{2}K_{\phi}(s,s_{0}))_{s\in\chi,s_{0}\in\chi_{0}}$ and $C_{00}=(\delta_{z}^{2}K_{\phi}(s_{0},s^{\prime}_{0}))_{s_{0},s_{0}^{\prime}\in\chi_{0}}$. Combined with the result in Proposition 1, the marginal predictive distribution for ${y}_{t0}$ is $t_{n_{t}}(\widetilde{X}_{0}{m}_{t,(1:p)}+C_{0}^{\mathrm{\scriptscriptstyle T}}{C}^{-1}{m}_{t,(p+1:p+n)},s_{t}(I_{n_{0}}+\widetilde{X}_{0}W_{t,(1:p,1:p)}\widetilde{X}_{0}^{\mathrm{\scriptscriptstyle T}}+C_{00}-C_{0}^{\mathrm{\scriptscriptstyle T}}{C}^{-1}C_{0}))$.
@@
Consider the assumptions in Proposition 1. The one-step ahead forecast distribution for the state vector and the corresponding one-step ahead predictive distribution are
${y}_{t+1}\mid{\theta}_{t+1},\sigma^{2},\mathcal{D}_{t}$ $\sim N\left(F_{t+1}{\theta}_{t+1},\ \sigma^{2}I_{n}\right)\;\mbox{ and }\;$
${\theta}_{t+1}\mid\sigma^{2},\mathcal{D}_{t}$ $\sim N\left(G_{t+1}{m}_{t},\ \sigma^{2}(G_{t+1}W_{t}G_{t+1}^{\mathrm{\scriptscriptstyle T}}+S_{t+1})\right)\;,$
respectively. The marginal predictive distribution for ${y}_{t+1}$ is $t_{n_{t}}(F_{t+1}G_{t+1}{m}_{t},s_{t}(I_{n}+F_{t+1}(G_{t+1}W_{t}G_{t+1}^{\mathrm{\scriptscriptstyle T}}+S_{t+1})F_{t+1}^{\mathrm{\scriptscriptstyle T}}))$. A general h-step ahead forecast can be obtained using recursive calculations.
@@ 5.2. Distribution theory for discrete time trajectory model
The following proposition provides the posterior predictive distributions for future points on a trajectory in the discrete time setup.
@@
Consider the setup leading to (3) and Proposition 4. Let $\Gamma_{0}=\{\gamma(1),\ldots,\gamma(T+1)\}$ be an enumeration of spatial locations and $\widetilde{\Gamma}_{0}=\{\widetilde{\gamma}_{1},\ldots,\widetilde{\gamma}_{n_{0}}\}\subseteq\Gamma_{0}$ be the set of $n_{0}$ distinct locations. Given a dataset $\mathcal{D}$ obtained up to time $T$, the posterior predictive distributions at time $T+1$ in (2) and (3) are
@@ 5.3. Distribution theory for continuous time trajectory model
The posterior predictive distributions for new points on a trajectory are obtained as follows.
@@
where $C_{z}=(\delta_{z}^{2}K_{\phi}((\gamma(t),t),(\gamma(t^{\prime}),t^{\prime})))_{t,t^{\prime}\in\mathcal{T}}$, $C_{z0}=(\delta_{z}^{2}K_{\phi}((\gamma(t),t),(\gamma(t_{0}),t_{0})))_{t\in\mathcal{T},t_{0}\in\mathcal{T}_{0}}$, $C_{z00}=(\delta_{z}^{2}K_{\phi}((\gamma(t_{0}),t_{0}),(\gamma(t_{0}^{\prime}),t_{0}^{\prime})))_{t_{0},t_{0}^{\prime}\in\mathcal{T}_{0}}$, $C_{\beta}=(\delta_{\beta}^{2}C_{\xi}(t,t^{\prime}))_{t,t^{\prime}\in\mathcal{T}}$, $C_{{\beta}0}=(\delta_{\beta}^{2}C_{\xi}(t,t_{0}))_{t\in\mathcal{T},t_{0}\in\mathcal{T}_{0}}$ and $C_{{\beta}00}=(\delta_{\beta}^{2}C_{\xi}(t_{0},t_{0}^{\prime}))_{t_{0},t_{0}^{\prime}\in\mathcal{T}_{0}}$. The marginal predictive distribution for ${y}_{0}$ is $t_{2a_{\sigma}^{*}}(\sum_{j=1}^{p}{x}_{j,0}C_{\beta 0}^{\mathrm{\scriptscriptstyle T}}{C}_{\beta}^{-1}\widehat{\beta}_{j}+C_{z0}^{\mathrm{\scriptscriptstyle T}}{C}_{z}^{-1}\widehat{z},(b_{\sigma}^{*}/a_{\sigma}^{*})(I_{n_{0}}+C_{z00}-C_{z0}^{\mathrm{\scriptscriptstyle T}}{C}_{z}^{-1}C_{z0}+\sum_{j=1}^{p}{x}_{j,0}(C_{\beta 00}-C_{\beta 0}^{\mathrm{\scriptscriptstyle T}}{C}_{\beta}^{-1}C_{\beta 0}){x}_{j,0}^{\mathrm{\scriptscriptstyle T}}))$, where $\widehat{z}$ and $\widehat{\beta}_{j}$ for $j=1,\ldots,p$ are the posterior means calculated as $m$ in Proposition 6.
5.4. Frequentist properties of posterior distributions
Theoretical investigations that shed light on the effectiveness of stacking in geostatistical settings have been explored by Zhang et al. (2023) in purely spatial contexts. Here, we investigate some theoretical results for the state-space setting. At the outset, it is worth recognizing that frequentist inference is rather limited because trajectory data, by definition, do not admit replicates at a single time point, while theoretical accessibility will require us to consider multiple, say $n$, spatial locations at each time point. The relevant setting here is the second adaptation of (1) discussed in Section 2.1, where we have multiple spatial locations at each epoch. For theoretical tractability, we consider spatial locations over Euclidean domains only.
For this development, we denote the response and the spatial process as $y_{t}(s)$ and $z_{t}(s)$, respectively, where $s$ is a generic spatial location in $\mathbb{R}^{d}$. We assume $n$ replicates of spatial locations $\chi_{n}=\{s_{1},\ldots,s_{n}\}$ at each $t$ and consider the model in (1) without the trend, i.e., $\beta_{t}=0$. Let $\mathcal{D}_{\chi_{n},t}$ now denote the entire dataset until time $t$ with spatial replicates in $\chi_{n}$. Hence,
@@ 5.4. Frequentist properties of posterior distributions
where $y_{t}$ and $z_{t}$ are each $n\times 1$ with elements $y_{t}(s_{i})$ and $z_{t}(s_{i})$, respectively, $\alpha$ is a fixed real number, and $K_{\phi}(s_{i},s_{j})=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{\|{s}_{i}-{s}_{j}\|}{\phi}\right)^{\nu}\mathcal{K}_{\nu}\left(\frac{\|{s}_{i}-{s}_{j}\|}{\phi}\right)$ is the Matérn kernel (Stein, 1999) defined for any pair of spatial locations ${s}_{i}$ and ${s}_{j}$ in a bounded region $\mathcal{S}\subset\mathbb{R}^{2}$. The parameters $\phi>0$ and $\nu>0$ model spatial decay and smoothness, respectively, and $\mathcal{K}_{\nu}$ is the modified Bessel function of the second kind of order $\nu$. Here, we fix $\nu$, call (14) Matérn model with parameters $\{\sigma,\phi,\delta_{z}\}$ and employ prior distributions $\sigma^{2}\sim IG(n_{\sigma}/2,n_{\sigma}s_{\sigma}/2)$ and ${z}_{0}\mid\sigma^{2}\sim N({m}_{0},\sigma^{2}S_{0})$.
@@
Lemma 1 implies that if the parameters are fixed at values different from the true (data generating) parameters, then the incorrectly specified model is equivalent to the model with the true parameters with regard to the distribution of $y$. This is an extension of Theorem 2.1 in Tang et al. (2021). Based on the fact, the following result on the error variance is available.
@@
where $a_{n}\asymp b_{n}$ means $a_{n}$ is bounded both above and below by $b_{n}$ asymptotically. If we set $m_{0}=0$, $S_{0}=K_{\phi^{\prime}}(\chi)$, $n_{\sigma}<\infty$, $s_{\sigma}<\infty$ and $\alpha<\infty$ in Matérn model (14), then the posterior distribution of $\sigma^{2}$ converges, as $n\to\infty$, to the degenerate distribution with entire mass at $\sigma^{2}_{\ast}$, i..e.,
@@
where $\rightsquigarrow$ denotes weak convergence of the probability measure and $\delta(x)$ is the Dirac measure at $x$.
Hence, if the spatial locations are not overly concentrated within $\mathcal{S}$, as the number of replicates increases, the posterior distribution of $\sigma^{2}$ degenerates to a point-mass distribution at the true parameter value (formally referred to as posterior strong consistency, Ghosal and van der Vaart, 2017). The assumption in Theorem 1 is explained in the Appendix in more detail.
@@
Under the assumptions in Theorem 1, the following two results hold for $t=1,\ldots,T$:
$\mathbb{E}_{*}[(Z_{tn}(s_{0})-\widetilde{z}_{t}(s_{0}))^{2}]=E_{n,t}^{A}+E_{n,t}^{B}~{}\mbox{ and }~{}\mathbb{E}_{*}\left[(Y_{tn}(s_{0})-\widetilde{y}_{t}(s_{0}))^{2}\right]-E_{n,t}^{A}-E_{n,t}^{B}\to 2\sigma_{*}^{2}\mbox{ as }n\to\infty\;,$
where $E_{n,t}^{A}=\sigma_{*}^{2}(\delta_{z}^{\prime 2}-\delta_{z}^{\prime 2}\{K_{\phi}(s,s_{0})\}_{s\in\chi_{n}}^{\mathrm{\scriptscriptstyle T}}K_{\phi}^{-1}(\chi_{n})\{K_{\phi}(s,s_{0})\}_{s\in\chi_{n}}+\{K_{\phi}(s,s_{0})\}_{s\in\chi_{n}}^{\mathrm{\scriptscriptstyle T}}K_{\phi}^{-1}(\chi_{n})R_{t}(R_{t}+I_{n})^{-2}R_{t}K_{\phi}^{-1}(\chi_{n})\{K_{\phi}(s,s_{0})\}_{s\in\chi_{n}}^{\mathrm{\scriptscriptstyle T}})+o(1)$, $E_{n,t}^{B}=\mathbb{E}_{*}[(\widetilde{z}_{t}(s_{0})-\{K_{\phi}(s,s_{0})\}_{s\in\chi}^{\mathrm{\scriptscriptstyle T}}K_{\phi}^{-1}(\chi){m}_{t})^{2}]$, $R_{t}=\alpha^{2}W_{t-1}+\delta_{z}^{{}^{\prime}2}K_{\phi^{\prime}}(\chi)$, $W_{t}=R_{t}-R_{t}^{\mathrm{\scriptscriptstyle T}}Q_{t}^{-1}R_{t}$ and $Q_{t}=R_{t}+I_{n}$.
The initial finding relates to the estimation of spatial-temporal effects. The estimation error at any given time is decomposed by $E_{n,t}^{A}$ and $E_{n,t}^{B}$. $E_{n,t}^{A}$ is the variance term, and its asymptotic form can be formulated without expectation. $E_{n,t}^{B}$ is a bias term, representing the difference between a true random variable and linear predictor by filtering $m_{t}$. The subsequent result indicates that the prediction error can be articulated in relation to the measurement error scale and the two terms.
As the number of measurement points increases while the observed area remains fixed, more points are close to the new points. Consequently, prediction accuracy is enhanced at new points, and $E_{n,t}^{A}$ and $E_{n,t}^{B}$ become small. An exploration of the convergence of $E_{n,t}^{B}$ in a limited scenario is presented in the Appendix. Additionally, because $E_{n,t}^{A}$ is challenging to analyze theoretically, we conducted a numerical examination of the decay of $E_{n,t}^{A}$ in the Appendix.
We develop the following result concerning stacking.
@@
Under the assumptions in Theorem 1, if $E_{n,t}^{B}\to 0$ as $n\to\infty$, it holds that
@@
This theorem ensures the validity of the stacking procedure. This result is not attributed to model averaging but stems from the asymptotic insignificance of parameter misspecification on the prediction model. However, remark that this is an asymptotic result, and for finite samples, the estimator of $z$ exhibits bias, and the prediction accuracy of $y$ is not stable. In this sense, model averaging is effective. Furthermore, in the context of statistical learning theory, it is known that while the convex hull increases the flexibility of the model (or reduces the training error), it does not increase its Rademacher complexity, i.e., the generalization gap. (See Chapter 6 in Mohri et al., 2018, for details). In other words, the stacked model has better predictive performance than a single model. For other theoretical justifications for stacking, see, e.g., van der Laan et al. (2007).

### v3

@@ 4. Prediction via stacking
We exploit the analytical closed forms for the posterior distributions and carry out inference using Bayesian stacking (Le and Clarke, 2017; Zhang et al., 2023). In both discrete-time and continuous-time trajectory models, we are able to obtain closed-form posterior distributions if we fix some hyperparameters in the spatial-temporal covariance structures. We consider a collection of $G$ models, $\{\mathcal{M}_{1},\ldots,\mathcal{M}_{G}\}$, where each $\mathcal{M}_{g}$ is specified by fixing a set of parameters so that the posterior distribution, given dataset $\mathcal{D}$, $p_{g}(\cdot\,|\,\mathcal{D})$, is in closed form.
Specifically for the discrete time model in (3), the posterior distribution for $\mathcal{M}_{g}$ is
@@
where $\delta^{2}_{g}=\{\delta^{2}_{\beta,g},\delta^{2}_{z,g}\}$, $\phi_{g}=\{\phi_{1,g},\phi_{2,g}\}$ and $\xi_{g}$ are the fixed values of these parameters for $\mathcal{M}_{g}$. The posterior predictive distributions of $y(\gamma(t_{0}),t_{0})$ and $z(\gamma(t_{0}),t_{0})$ at the new time point $t_{0}$ are calculated from t-distributions; see Section 5.3 for details.
@@ 4.2. Predictive stacking of distributions
Spatial energetics pursues predictive inference on trajectories entailing interpolation of the latent process at arbitrary points. This drives predictions for the outcomes. We achieve this by stacking the posterior predictive distributions for each $\mathcal{M}_{g}$ using $\mathcal{D}_{train}$, which is a multivariate t-distribution. Similar to the stacking of means, we consider the weights. Stacking maximizes the score function $S\left(\sum_{g=1}^{G}a_{g}p_{g}(\cdot\mid\mathcal{D}_{train}),q_{t}(\cdot\mid\mathcal{D}_{train})\right)$ to obtain the weights, where $q_{t}$ is a posterior distribution with true underlying parameters. If we employ a logarithmic score, corresponding to the Kullback–Leibler divergence (Yao et al., 2018; Zhang et al., 2023), the weights are obtained as $\operatornamewithlimits{argmax}_{a_{1},\ldots,a_{G}}\sum_{y\in\mathcal{D}_{valid}}\log\left(\sum_{g=1}^{G}a_{g}p_{g}(y\mid\mathcal{D}_{train})\right)$, the logarithm acts on the pseudo-posterior probabilities, given the weighted models. Thus, we define the distributional prediction by maximizing the pseudo-log joint posterior probability. Note that $p_{g}(\cdot\mid\mathcal{D}_{train})$ is a multivariate t-distribution, and hence, evaluating the posterior probability of the validation data is readily available. This optimization problem can be solved as a linearly constrained problem via an adaptive barrier algorithm (Lange, 2010). Algorithm 2 presents the steps involved in stacking of predictive densities.
@@
The $G$ candidate models in Algorithms 1 and 2 are computed in parallel and computation of the weights is negligibly small compared to that of the posterior distribution. Further, optimization is supported by many packages in various statistical programming languages. In particular, for the subsequent illustrations, we employed the (“stats” and “quadprg” packages R Core Team, 2024; Turlach and Weingessel, 2019) in the R statistical computing environment. By contrast, MCMC demands a substantial number of iterations for convergence, and the issue is exacerbated with the larger values of $n$ and $T$.
@@ 5.1. Distribution theory for Bayesian DLMs
Fixing $\delta_{\beta},\delta_{z},\phi$ yields closed-form posterior distributions for ${\theta}_{t}$ and $\sigma$, which aids stacking. We collect recursion equations used in calculating the posterior distribution for (1).
@@
where ${m}_{t,(1:p)},{m}_{t,(p+1:p+n)}$ are the first $p$ elements and the remaining elements of ${m}_{t}$, $W_{t,(1:p,1:p)}$ is the top-left $p\times p$ square of $W_{t}$, $C=(\delta_{z}^{2}K_{\phi}(s,s^{\prime}))_{s,s^{\prime}\in\chi}$, $C_{0}=(\delta_{z}^{2}K_{\phi}(s,s_{0}))_{s\in\chi,s_{0}\in\chi_{0}}$ and $C_{00}=(\delta_{z}^{2}K_{\phi}(s_{0},s^{\prime}_{0}))_{s_{0},s_{0}^{\prime}\in\chi_{0}}$. Combined with Proposition 1, the marginal predictive distribution for ${y}_{t0}$ is $t_{n_{t}}(\widetilde{X}_{0}{m}_{t,(1:p)}+C_{0}^{\mathrm{\scriptscriptstyle T}}{C}^{-1}{m}_{t,(p+1:p+n)},s_{t}(I_{n_{0}}+\widetilde{X}_{0}W_{t,(1:p,1:p)}\widetilde{X}_{0}^{\mathrm{\scriptscriptstyle T}}+C_{00}-C_{0}^{\mathrm{\scriptscriptstyle T}}{C}^{-1}C_{0}))$.
@@
Consider the assumptions in Proposition 1. The one-step ahead forecast and corresponding predictive distributions for the state vector are ${y}_{t+1}\mid{\theta}_{t+1},\sigma^{2},\mathcal{D}_{t}\sim N\left(F_{t+1}{\theta}_{t+1},\ \sigma^{2}I_{n}\right)$ and ${\theta}_{t+1}\mid\sigma^{2},\mathcal{D}_{t}\sim N\left(G_{t+1}{m}_{t},\ \sigma^{2}(G_{t+1}W_{t}G_{t+1}^{\mathrm{\scriptscriptstyle T}}+S_{t+1})\right)$, respectively. The predictive distribution for ${y}_{t+1}$ is $t_{n_{t}}(F_{t+1}G_{t+1}{m}_{t},s_{t}(I_{n}+F_{t+1}(G_{t+1}W_{t}G_{t+1}^{\mathrm{\scriptscriptstyle T}}+S_{t+1})F_{t+1}^{\mathrm{\scriptscriptstyle T}}))$. A general h-step ahead forecast can be obtained using recursive calculations.
@@ 5.2. Distribution theory for discrete time trajectory model
The posterior predictive distributions for future points on a trajectory is also available.
@@
Consider the setup for (3) and Proposition 4. Let $\Gamma_{0}=\{\gamma(1),\ldots,\gamma(T+1)\}$ be an enumeration of spatial locations and $\widetilde{\Gamma}_{0}=\{\widetilde{\gamma}_{1},\ldots,\widetilde{\gamma}_{n_{0}}\}\subseteq\Gamma_{0}$ be the set of $n_{0}$ distinct locations. Given $\mathcal{D}$ up to time $T$, the predictive distributions at $T+1$ in (2) and (3) are
@@ 5.3. Distribution theory for continuous time trajectory model
The predictive distributions for new points on a trajectory are obtained as follows.
@@
where $C_{z}=(\delta_{z}^{2}K_{\phi}((\gamma(t),t),(\gamma(t^{\prime}),t^{\prime})))_{t,t^{\prime}\in\mathcal{T}}$, $C_{z0}=(\delta_{z}^{2}K_{\phi}((\gamma(t),t),(\gamma(t_{0}),t_{0})))_{t\in\mathcal{T},t_{0}\in\mathcal{T}_{0}}$, $C_{z00}=(\delta_{z}^{2}K_{\phi}((\gamma(t_{0}),t_{0}),(\gamma(t_{0}^{\prime}),t_{0}^{\prime})))_{t_{0},t_{0}^{\prime}\in\mathcal{T}_{0}}$, $C_{\beta}=(\delta_{\beta}^{2}C_{\xi}(t,t^{\prime}))_{t,t^{\prime}\in\mathcal{T}}$, $C_{{\beta}0}=(\delta_{\beta}^{2}C_{\xi}(t,t_{0}))_{t\in\mathcal{T},t_{0}\in\mathcal{T}_{0}}$ and $C_{{\beta}00}=(\delta_{\beta}^{2}C_{\xi}(t_{0},t_{0}^{\prime}))_{t_{0},t_{0}^{\prime}\in\mathcal{T}_{0}}$. The predictive distribution for ${y}_{0}$ is $t_{2a_{\sigma}^{*}}(\sum_{j=1}^{p}{x}_{j,0}C_{\beta 0}^{\mathrm{\scriptscriptstyle T}}{C}_{\beta}^{-1}\widehat{\beta}_{j}+C_{z0}^{\mathrm{\scriptscriptstyle T}}{C}_{z}^{-1}\widehat{z},(b_{\sigma}^{*}/a_{\sigma}^{*})(I_{n_{0}}+C_{z00}-C_{z0}^{\mathrm{\scriptscriptstyle T}}{C}_{z}^{-1}C_{z0}+\sum_{j=1}^{p}{x}_{j,0}(C_{\beta 00}-C_{\beta 0}^{\mathrm{\scriptscriptstyle T}}{C}_{\beta}^{-1}C_{\beta 0}){x}_{j,0}^{\mathrm{\scriptscriptstyle T}}))$, where $\widehat{z}$ and $\widehat{\beta}_{j}$ for $j=1,\ldots,p$ are the posterior means calculated as $m$ in Proposition 6.
5.4. Repeated sampling properties of posterior distributions
Here, we investigate some theoretical results for the state-space setting. At the outset, it is worth recognizing that classical inference is rather limited because trajectory data, by definition, do not admit replicates at a single time point. Theoretical accessibility requires multiple, say $n$, spatial locations at each time point. The relevant setting here is the second adaptation of (1) discussed in Section 2.1 with multiple spatial locations at each epoch.
For convenient notation, let $y_{t}(s)$ and $z_{t}(s)$ denote the response and the spatial process, respectively, where $s$ is a generic spatial location in $\mathbb{R}^{d}$. We assume $n$ spatial replicates $\chi_{n}=\{s_{1},\ldots,s_{n}\}$ at each $t$ and consider (1) without the trend, i.e., $\beta_{t}=0$. Let $\mathcal{D}_{\chi_{n},t}$ now be the entire dataset until time $t$ with spatial replicates in $\chi_{n}$. Hence,
@@ 5.4. Repeated sampling properties of posterior distributions
where $y_{t}$ and $z_{t}$ are each $n\times 1$ with elements $y_{t}(s_{i})$ and $z_{t}(s_{i})$, respectively, $\alpha$ is a fixed real number, and $K_{\phi}(s_{i},s_{j})=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{\|{s}_{i}-{s}_{j}\|}{\phi}\right)^{\nu}\mathcal{K}_{\nu}\left(\frac{\|{s}_{i}-{s}_{j}\|}{\phi}\right)$ is the Matérn kernel (Stein, 1999) defined for any pair of spatial locations ${s}_{i}$ and ${s}_{j}$ in a bounded region $\mathcal{S}\subset\mathbb{R}^{2}$. The parameters $\phi>0$ and $\nu>0$ model spatial decay and smoothness, respectively, and $\mathcal{K}_{\nu}$ is the modified Bessel function of the second kind of order $\nu$. Here, we fix $\nu$and employ prior distributions $\sigma^{2}\sim IG(n_{\sigma}/2,n_{\sigma}s_{\sigma}/2)$ and ${z}_{0}\mid\sigma^{2}\sim N({m}_{0},\sigma^{2}S_{0})$.
@@
Lemma 1 implies that if the parameters are fixed at values different from the true (data generating) parameters, then the incorrectly specified model is equivalent to that with the true parameters with regard to the distribution of $y$. This extends Theorem 2.1 in Tang et al. (2021). Based on this, the following result on the error variance is available.
@@
where $a_{n}\asymp b_{n}$ means $a_{n}$ is bounded both above and below by $b_{n}$ asymptotically. If we set $m_{0}=0$, $S_{0}=K_{\phi^{\prime}}(\chi)$, $n_{\sigma}<\infty$, $s_{\sigma}<\infty$ and $\alpha<\infty$ in the Matérn model (14), then the posterior distribution of $\sigma^{2}$ converges, as $n\to\infty$, to the degenerate distribution at $\sigma^{2}_{\ast}$, i.e.,
@@
where $\rightsquigarrow$ denotes weak convergence of $p(\sigma^{2}\mid\mathcal{D}_{\chi_{n},t})$ and $\delta(x)$ is Dirac measure at $x$.
If the spatial locations are not overly concentrated within $\mathcal{S}$, then, as the number of replicates increases, the posterior distribution of $\sigma^{2}$ degenerates to a point-mass distribution at the true parameter value (formally referred to as posterior strong consistency, Ghosal and van der Vaart, 2017). See the Appendix for details on the assumption in Theorem 1.
@@
With assumptions in Theorem 1, the following results hold for $t=1,\ldots,T$:
$\mathbb{E}_{*}[(Z_{tn}(s_{0})-\widetilde{z}_{t}(s_{0}))^{2}]=E_{n,t}^{A}+E_{n,t}^{B}~{}\mbox{ and }~{}\mathbb{E}_{*}\left[(Y_{tn}(s_{0})-\widetilde{y}_{t}(s_{0}))^{2}\right]-E_{n,t}^{A}-E_{n,t}^{B}\to 2\sigma_{*}^{2}~{}\mathrm{as}~{}n\to\infty,$
where $E_{n,t}^{A}=\sigma_{*}^{2}\left(\delta_{z}^{\prime 2}h_{\phi}(\chi_{n},s_{0})+g(\chi_{n},s_{0})\right)+o(1)$, $h_{\phi}(\chi_{n},s_{0})$ and $g_{\phi}(\chi_{n},s_{0})$ are functions defined as $h_{\phi}(\chi_{n},s_{0})=1-K_{\phi}(\chi_{n},s_{0})^{\mathrm{\scriptscriptstyle T}}K_{\phi}^{-1}(\chi_{n})K_{\phi}(\chi_{n},s_{0})$ and $g(\chi_{n},s_{0})=K_{\phi}(\chi_{n},s_{0})^{\mathrm{\scriptscriptstyle T}}K_{\phi}^{-1}(\chi_{n})R_{t,n}(R_{t,n}+I_{n})^{-2}R_{t,n}K_{\phi}^{-1}(\chi_{n})K_{\phi}(\chi_{n},s_{0})^{\mathrm{\scriptscriptstyle T}}$, $K_{\phi}(\chi_{n},s_{0})=\{K_{\phi}(s,s_{0})\}_{s\in\chi_{n}}$ is $n\times 1$, and $E_{n,t}^{B}=\mathbb{E}_{*}[(\widetilde{z}_{t}(s_{0})-K_{\phi}(\chi_{n},s_{0})^{\mathrm{\scriptscriptstyle T}}K_{\phi}^{-1}(\chi_{n}){m}_{t})^{2}]$ with $R_{t,n}=\alpha^{2}W_{t-1}+\delta_{z}^{{}^{\prime}2}K_{\phi^{\prime}}(\chi_{n})$, $W_{t}=R_{t,n}-R_{t,n}^{\mathrm{\scriptscriptstyle T}}Q_{t,n}^{-1}R_{t,n}$ and $Q_{t,n}=R_{t,n}+I_{n}$.
The initial finding relates to estimates of spatial-temporal effects. The estimation error at time $t$ is decomposed by $E_{n,t}^{A}$ and $E_{n,t}^{B}$, where $E_{n,t}^{A}$ is the variance term with its asymptotic form expressible without expectations, and $E_{n,t}^{B}$ is a bias term representing the difference between a true random variable and linear predictor by filtering $m_{t}$. The subsequent result indicates that the prediction error is articulated in relation to the measurement error scale and the two terms. As the number of measurement points increases while the observed area remains fixed, more points are close to the new points. Consequently, prediction accuracy is enhanced at new points, and $E_{n,t}^{A}$ and $E_{n,t}^{B}$ become small. An exploration of the convergence of $E_{n,t}^{B}$ in a limited scenario is presented in the Appendix. Since $E_{n,t}^{A}$ is challenging to analyze theoretically, we present numerical experiments on the decay of $E_{n,t}^{A}$ in the Appendix. We develop the following result concerning stacking.
@@
Under the assumptions in Theorem 1, if $E_{n,t}^{B}\to 0$ as $n\to\infty$, then
@@
This theorem validates stacking as an inferential procedure. This is an asymptotic result so for finite samples the estimator of the process exhibits bias which can diminish predictive accuracy. Model averaging enhances stability and, in the context of statistical learning, the convex hull increases the flexibility of the model (or reduces the training error) without increasing its Rademacher complexity, i.e., the generalization gap. (See Chapter 6 in Mohri et al., 2018, for details). Hence, stacked inference delivers better predictive performance than a single model (see van der Laan et al., 2007, for further theoretical results).